OWG Meeting 10-Sep-2014
Location and Time
Room: CC F326-327
(if problems, call phone in conference room: 757-269-6460)
- To join via Polycom room system go to the IP Address: 22.214.171.124 (bjn.vc) and enter the meeting ID: 120390084.
- To join via a Web Browser, go to the page https://bluejeans.com/120390084.
- To join via phone, use one of the following numbers and the Conference ID: 120390084
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- More information on connecting to bluejeans is available.
*note: you'll be muted when first connected to the meeting. press *4 to unmute
- Collaboration Meeting
- Major systems reports
- DAQ Config. Parameters in Data Stream
- Mantis Task Tracker
Attendees: Alex S., Bryan M., David L., Simon T., Sergey F., Curtis M., Hovanes E., Dave A., Sergeui P., Beni Z.
- There was a power outage in the counting house today. The power is now back on. Cause is unknown. It is also unknown which if any gluon machines were affected since they should mostly be on UPS.
- Reinstalled 22 CTP boards after replacing thermal pads
- Still have some funny issues with temperature monitoring on a couple of the FPGAs
- Chips did not seem too warm to the touch so we don't think anything is being damaged
- Currently have 2 algorithms running on GTP
- BCAL north cosmic trigger has been tested and is working
- Ran at ~200Hz for >20hrs
- Next step is to get full BCAL trigger working. After that, TD's will all be moved into a single TS crate
- Trigger commissioning plan is under development
- Will come after the photon beam is established in the overall commissioning plan.
- Tagger microscope will be moving into Tagger Hall next Friday.
- Do not plan to have tagger electronics in hall during commissioning.
- Still waiting to test full system (need all TD's in single TS crate)
- Currently expect to run up to 8kHz and testing so far indicates this should not be a problem
- Preparing OpenSUSE node (gluon49) to test newest Linux kernel for DAQ system crashes
- This will likely be done during the large scale DAQ system testing tomorrow
- CODA 3.02 is n final testing stages. Could be ready for use in Hall-D as early as tomorrow
- Will also be installed in /site once it is ready
- Potential performance improvements are likely to push up the current 8kHz limit
- Several bug fixes including the one that causes crashes in the secondary ROL at end of run
- A graphical monitoring program "daqmon" has been developed
- Presents histograms generated on ROCs
- Uses dedicated communication mechanism in addition to RootSpy's cMsg based system
- Allows for faster updates for some complex displays where greater than 1Hz update rate is desired
- Start.stop scripts are now ready and need testing with DAQ system
- Starts/Kills monitoring processes on multiple remote computers
- Supports multiple monitoring levels (occupancy, reconstruction,...)
- GUI has been built that presents integrated rates and lists active nodes.
- Can be used to start and stop monitoring processes, but usually that should be done by DAQ system
- Vanik and Nerses finished with the Voltage Control Apps.
- Not all of them have been tested by the detector groups yet
- Next focus is on the motor apps which will include the harp scan software
- Work has started on readout of 2 Beam Position Monitors
- First attempt tried to port existing software from VxWorks to Linux. This proved more complex than first anticipated
- Reverted to using VxWorks where driver for Struck VME scalers already exist.
- Hovanes has been working with Mark Ito on the alarm system.
- Alarm system configuration can be dumped into an XML file and read back in
DAQ Config. Parameters in Data Stream
- David presented a scheme for implementing some DAQ configuration parameters into the data stream
- System defines an EVIO bank that will be written out by each crate for each DAQ event
- Initial estimate is roughly 1kB per DAQ event or about a 5% increase for a 19.4kB event size
- If in multi-event block mode, percentage decreases (should be O(0.1%) for high rate running)